# JWST Discovery of a High-Redshift Tidal Disruption Event Candidate in COSMOS-Web

### v1

@@ 3.2 Search Results
We ultimately selected 1764 sources that passed our point source cuts, 117 passed our color-color, and magnitude cuts. We visualize these candidates in color-color and color-magnitude space in Figure 5. Among these sources, 82 of them had no listing in the COSMOS2020 catalog. We divide these into “bright” and “faint” sources. “Bright” sources are sources with $M_{F115W}>27.8$ AB mag, where $27.8$ mag is the detection limit in HST F814W in COSMOS2020, the deepest imaging of this field. These filters do not overlap– the $\lambda_{\rm max}$ of F814W is $9643$ and the $\lambda_{\rm min}$ of F115W is $9975$– but they are near enough to where nondetection in F115W approximates a nondetection in F814W. Therefore, “bright” sources should have mostly been detected in HST. We find that our candidates contain 79 “bright” sources and 3 “faint sources”. For all 79 sources without HST catalog entries, we download the HST F814W cutouts, and search through them by hand for nondetections. We ultimately determine by visual inspection that 72 of the 79 “bright” sources which pass our cuts do have a counterpart in F814W images, but are excluded from COSMOS2020. We note that these sources often lie on the edge of a tile, in the masked region near a bright star, or have inaccurate astrometry. Among the sources not seen in the manual inspection, six are very red and not bright in F184W. We are left with one source which is detected in COSMOS-Web but not in COSMOS2020. The remaining candidate peaks at $\sim 25$ mag in F150W, which is $>1$ mag above the limit of previous infrared imaging with UltraVISTA. This source is not detected in UVISTA data to even $1\sigma$, confirming its nature as a transient and viable high-redshift TDE candidate. We will refer to this TDE candidate at HZTDE-1.
@@
Figure 6: Top: Example of a high-redshift TDE candidate which we believe to be a $z=3$ galaxy contaminant. Despite having a blackbody-like SED shape in NIRCam, the HSC photometry shows that the SED is flat in the optical, with a Lyman break beginning in the g-band. Its F770W fluc greatly exceeds that of a blackbody. The source appears point-like in F115W and F150W, but in the F277W imaging it can be seen to be slightly extended, although compact. The best-fit galaxy template is plotted in red, star template in light blue, and quasar in purple. Bottom: Example of a high-redshift TDE candidate which we believe to be a local cool dwarf star. The source has no listing in the Hubble Source catalog, although it is detected in COSMOS2020. This source has no Lyman break, and is detected to $5\sigma$ in every HSC band.
@@
These candidates are selected only by JWST photometry, and the HSC photometry is consistent with a blackbody with the same peak as a high redshift TDE for the vast majority of sources, as seen in Figures 6 and 7. Figure 7 shows an example fit of our simplified high-redshift TDE model to HZTDE-1, characterized only by absolute magnitude, blackbody temperature, and redshift, to the photometry. We find that, while the model fits well to the majority of our data, the degeneracy between redshift and temperature means that the fits are extremely dependent on the starting guess and there is no way to discern a source’s redshift without observing a Lyman dropout. Additionally, shown as a comparison, this could also match the SED of an M-dwarf star. Therefore, for these data it is essential that we utilize all previous HST imaging to check for variability and optical HSC data to look for a dropout.
@@
Objects without Lyman dropout are likely cool dwarf stars. Stars with temperatures $2400\textrm{K}<T<3500\textrm{K}$ have blackbody peaks around 1.2 $\mu$m, the same observed wavelength as a $z>5$ TDE. These stars are abundant in the Milky Way, although are less common when looking out of the Galactic Plane. Figure 6 shows a source which we believe is a star. It has no listing in the COSMOS2020 catalog, but manual inspection can identify the corresponding object in previous COSMOS imaging.
@@ 4 Properties of the TDE Candidate HZTDE-1
Figure 7: Cutouts (top panels) and SED (bottom panel) of high redshift TDE candidate, HZTDE-1. The cutouts include imaging from UltraVISTA, Hyper Suprime Cam, Hubble ACS, and JWST NIRCam. The source is only detected to $>3\sigma$ in the NIRCam images. The SED shows the observed SED as red circles, synthetic photometry for the best-fit galaxy spectrum (purple line) and TDE spectrum (red line) in diamonds, and all photometric upper limits as downward-pointing triangles. Errorbars in magnitude exist for the NIRCam data but are too small to be seen. The transmission curve for each filter is plotted below it. A simulated spectrum of TDE 2019qiz at $z=5$ is overplotted as a dotted line, and a template for a $24,000$ K superluminous supernova at $z=3.6$ is the yellow line.
@@ 4.1 Comparison to supernova models
There are three nearby galaxies that HZTDE-1 could belong to if it is a lower-redshift supernova. We assess its possible association with the galaxy directly above it in the cutouts shown in Figure 7, because it is the closest galaxy, with an angular separation of $0.998$”, the brightest galaxy with $M_{F444W}=23.7$ AB mag, and the most extended with effective radius $R_{e}=0.20$” as modeled by GALFIT (Peng et al., 2002). The GALFIT model for the sources in the field is shown in Figure 9, alongside the F [^4] W image and the residual. The best-fit galaxy spectrum to this source’s SED has a photometric redshift $z_{\rm phot}=1.75^{+0.1}_{-0.07}$. The SED is well-sampled from the HSC $g$-band to NIRCam F [^4] W, and the model spectrum fits accurately with a $\chi^{2}=1.72$ per degree of freedom. We follow the technique developed in Gupta et al. (2016) to assess host galaxy association for a given supernova through calculating the directional light radius (DLR), which is the elliptical radius of a galaxy in the direction of the supernova. They then calculate the $d_{\rm DLR}$, which is the distance between the supernova and the galaxy normalized by the DLR. They utilize both $HST$ ACS data and MICECATv2.0 simulations to show that any SN with a galaxy $>5$ DLR away should be assigned “hostless”, and achieve $\sim 95\%$ completeness with a similar cut. We calculate the unitless dDLR for HZTDE-1 to be $5.05$ relative to the nearest galaxy, making it highly likely to be a hostless transient. This dDLR is even higher for the short-wavelength NIRCam filters ($>6$ DLR in F115W), which have higher spatial resolution.
Figure 9: GALFIT modeling of the morphology of the sources in the F [^4] W imaging of HZTDE-1. We display the size of $1$ directional light radius (DLR) from the galaxy to the transient to show the accuracy of the modeling. The HZTDE-1 is $>5$ DLR from the galaxy, unlikely to be associated. In shorter-wavelength filters, this DLR separation increases.
We assume that HZTDE-1 is unlikely to be a “hostless” non-superluminous supernova. However, given the rare case that HZTDE-1 is associated with the nearest galaxy despite extreme distance, we fit supernova SEDs for each supernova subtype at its photometric redshift, $z=1.75\pm 0.10$. We use the full range of supernova properties described in Section 2.5. For every supernova subtype, we fit every available model in sncosmo (Barbary et al., 2016), extending to the infrared with a blackbody when the model has insufficient wavelength range. We do not fit models for Population III SNe, Pair-Instability Supernovae (PISNe), or other exotic transients. We show our best-fit supernova spectra in Figure 10, and their respective lightcurves in Figure 11. We find that HZTDE-1 is significantly too red to be a SN Ia at $z\sim 1.75$ and too bright to be a SN Ia at $z>2$. HZTDE-1 is $\gtrapprox 1$ too bright to be explained by a Ib/c at $z\sim 1.75$. The SED is somewhat plausibly explained by an anomalously bright IIP before maximum brightness or a bright IIn prior to maximum brightness, but both under-predict the flux in F [^4] W at $z\sim 1.75$, which would be $1.6~{}\mu$m in the rest-frame. These best models are shown in Figure 10. This excess flux could be explained by contribution from thermal dust emission, but given the extreme offset from the potential host it is unlikely that the supernova would be embedded in dust. In summary, for HZTDE-1 to be a supernova associated with its nearest galaxy it must be an anomalously bright SN IIn at an extreme offset with an infrared flux excess.
@@
There is a possibility that HZTDE-1 is a superluminous supernova (SLSN), as shown in Figure 7. SLSNe have similar intrinsic luminosities to TDEs and their spectral continua are primarily blackbody thermal emission (Gomez et al., 2024). Their host galaxies tend to be less massive, meaning that at large enough distances ($z\gtrsim 3$) they may also outshine their hosts and be point sources. Therefore, to be certain whether HZTDE-1 is a high-redshift TDE or high-redshift hostless supernova, either follow-up spectroscopy or monitoring of its photometric evolution is necessary. With a single photometric detection, the relative rates of TDEs and SLSNe at $z>3$ is the main factor that determines what HZTDE-1 likely is. Current best estimates of the volumetric rate of SLSNe place them somewhere between $0.91\times 10^{-7}~{}\rm{SNe}/\rm{year}/\rm{Mpc}^{3}$ at $z=1$ (Prajs et al., 2017) and $1.99\times 10^{-7}~{}\rm{SNe}/\rm{year}/\rm{Mpc}^{3}$ at $z=0.2$ (Quimby et al., 2013), as compared to the current best estimate for the local optical TDE volumetric rate of $3.1\times 10^{-7}~{}\rm{TDEs}/\rm{year}/\rm{Mpc}^{3}$. Tanaka et al. (2013) predicts the volumetric rate at SLSNe at higher redshifts (up to $z\geq 10$) by scaling it with the cosmic star formation rate (SFR) density using two SFR models (Robertson & Ellis, 2012; Hopkins & Beacom, 2006). They calibrate to the SLSN low-redshift rate, and account for the initial mass function and the fraction of massive stars to find that the SLSN volumetric rate increases with redshift until $z\sim 5$, where the SFR density peaks, and then the rate decreases with redshift indefinitely. At its maximum, $z\sim 5$, the volumetric SLSN rate is 10×10$\times$10 × the local rate. We expect the TDE rate to also be elevated by at least 10×10$\times$10 × the local rate, as shown in Karmen & et al. (in prep.). HZTDE-1 would need to be extremely blue for a SLSN, with a best-fit temperature $24,000$ K at best-fit redshift $z=3.6$, but this is possible briefly before peak and during peak brightness. While this makes a SLSN somewhat unlikely, given the comparable rates there is no way to rule out a SLSN without further photometric monitoring or spectroscopy.
Figure 11: Lightcurves of best-fit supernovae models to HZTDE-1, given that they are at redshift within $0.1$ of the closest galaxy. Each panel is in a single COSMOS-Web NIRCam filter. The best-fitting models, IIP and IIn, do not adequately explain the F [^4] W flux, which would be at $\sim 1.6\mu$m in the rest frame.
@@ 5 Discussion and Future Confirmation
In this work, we model the properties of TDEs at high redshift assuming that their populations are similar to those at low redshift. We argue that the assumptions made about their host galaxies (Section 2.1) and their colors (Section 2.3) are treated as bounds rather than constraints and that evolution of these properties with redshift will lie within our selection criteria. However, any number of these assumptions can be relaxed. In future work, for example, modeling of TDEs in extended hosts, or modeling of dust-extincted TDE emission can increase the efficiency of searches for TDEs in wide-field surveys. As this search offers an opportunity to study the UV emission shortward of $2000$ which is usually not observed, deviations from blackbody assumptions about TDE spectra, such as those mentioned in Section 2.4 may emerge from studies of high-redshift TDEs.
Both spectroscopic and photometric follow-up is needed to confirm HZTDE-1’s classification as a TDE and its redshift. TDEs have blue spectra with blackbody continua and broad Hydrogen and/or Helium emission features that narrow as the TDE evolves. A key distinction between TDEs and supernovae is their decline: TDEs fade slowly, typically as $t^{-5/3}$, and do not change temperature as they cool (van Velzen et al., 2020). Without imaging of a host galaxy to verify that HZTDE-1 is nuclear, combining identification these spectroscopic features and photometric monitoring is needed to confidently confirm it as a TDE. Spectroscopy, specifically in the infrared with JWST’s NIRSpec, is key to observing key TDE emission lines. X-ray emission is not expected because, as mentioned, the CGM will absorb all of the soft X-ray emission characteristic of TDEs at $z>3$. Longer wavelengths may be worth monitoring for later-time infrared flares due to dust echoes (van Velzen et al., 2016).
@@ 5.1 TDE rates in COSMOS-Web
In Karmen & et al. (in prep.), we calculate observed TDE rates as a function of redshift in LSST, Roman, and the COSMOS-Web survey. We find that if the TDE rate does not evolve as a function of redshift, there is a $68\%$ chance of observing a single TDE in COSMOS-Web, and a $7\%$ chance of observing a TDE at $z>4$. We show, however, that the observed properties of high-redshift galaxies would imply a dramatically enhanced rate of TDEs at high-redshift $105<z<$ 10. Namely, many galaxies are observed to be centrally compact (Finkelstein et al., 2023; Kartaltepe et al., 2023; Baggen et al., 2023; Guia et al., 2024), and may host nuclear stellar clusters (Ricotti et al., 2016; Sun et al., 2023) at high redshifts. This increases central stellar density, which in turn will lead to a ×$\times 100$ TDE rate enhancement (Kritos et al., 2024; Pfister et al., 2020). Additionally, galaxy mergers are shown to enhance TDE rates both observationally and theoretically. Host studies of optical/UV TDEs demonstrate a clear preference for galaxies which are believed to be products of mergers (French et al., 2020; Wevers & French, 2024), and the TDE rate is seen in hydrodynamical simulations to be enhanced by an order of magnitude for 100 Myr after a merger (Pfister et al., 2021). With a galaxy merger rate that increases as a function of redshift, peaking at cosmic noon (Ventou et al., 2017; Duncan et al., 2019; Perna et al., 2025) we predict TDE rates to be enhanced from mergers as well. In the past it was expected that TDE rates decrease with the SMBH mass function as a function of redshift. However, given the abundance of supermassive and overmassive black holes detected at early cosmic times (e.g. Maiolino et al., 2023; Larson et al., 2023; Kokorev et al., 2023), this phenomenon will be significantly reduced. TDEs are among the few luminous phenomena very close to an SMBH, and are believed to produce super-Eddington accretion at early times (Strubbe & Quataert, 2009) so their study at high redshifts can give an indication to the mechanisms by which early SMBHs accrete and grow.

### v4

@@ 3.2 Search Results
We ultimately selected 1764 sources that passed our point source cuts, 117 passed our color-color, and magnitude cuts. We select sources based on COSMOS-Web NIRCam colors alone, in order to prevent time evolution between different surveys from affecting calculated colors. We visualize these candidates in color-color and color-magnitude space in Figure 6.
@@
Among these sources, 82 of them had no listing in the COSMOS2020 catalog. We divide these into “bright” and “faint” sources. “Bright” sources are sources with $M_{F115W}>27.8$ AB mag, where $27.8$ mag is the detection limit in HST F814W in COSMOS2020, the deepest imaging of this field. These filters do not overlap– the $\lambda_{\rm max}$ of F814W is $9643$ Å and the $\lambda_{\rm min}$ of F115W is $9975$ Å– but they are near enough to where nondetection in F115W approximates a nondetection in F814W. Therefore, “bright” sources should have mostly been detected in HST. We find that our candidates contain 79 “bright” sources and 3 “faint sources”. For all 79 sources without HST catalog entries, we download the HST F814W cutouts, and search through them by hand for nondetections. We ultimately determine by visual inspection that 72 of the 79 “bright” sources which pass our cuts do indeed have a counterpart in F814W images, but are excluded from COSMOS2020. We note that these sources often lie on the edge of a tile, in the masked region near a bright star, or have inaccurate astrometry. Among the sources not seen in F814W in manual inspection, six of them appear to be static, yet are too faint in F814W to be detected by previous HST imaging. We are left with one source which is detected in COSMOS-Web but not in COSMOS2020. The remaining candidate peaks at $\sim 25$ mag in F150W, which is >1>1 mag above the limit of previous infrared imaging with UltraVISTA. This source is not detected in UVISTA data to even $1\sigma$, confirming its nature as a transient and viable high-redshift TDE candidate. We will refer to this TDE candidate at HZTDE-1.
@@
Figure 7: Top: Example of a high-redshift TDE candidate which we believe to be a $z=3$ galaxy contaminant. Despite having a blackbody-like SED shape in NIRCam, the HSC photometry shows that the SED is flat in the optical, with a Lyman break beginning in the g-band. Its F770W flux greatly exceeds that of a blackbody. The source appears point-like in F115W and F150W, but in the F277W imaging it can be seen to be slightly extended, although compact. The best-fit galaxy template is plotted in red, star template in light blue, and quasar in purple. This galaxy does not have a F814W flux measurement in the COSMOS2020 catalog due to photometry flags, but manual inspection confirms it was detected. Bottom: Example of a high-redshift TDE candidate which we believe to be a local cool dwarf star. The source has no listing in the Hubble Source Catalog, although it is detected in COSMOS2020. This source has no Lyman break, and is detected to $5\sigma$ in every HSC band.
@@
These candidates are selected only by JWST photometry, and the HSC photometry is consistent with a blackbody with the same peak as a high redshift TDE for the vast majority of sources, as seen in Figures 7 and 8. Figure 8 shows an example fit of our simplified high-redshift TDE model to HZTDE-1, characterized only by absolute magnitude, blackbody temperature, and redshift, to the photometry. We find that, while the model fits well to the majority of our data, the degeneracy between redshift and temperature means that the fits are extremely dependent on the starting guess and there is no way to discern a source’s redshift without observing a Lyman dropout. Additionally, shown as a comparison, this could also match the SED of an M-dwarf star. Therefore, for these data it is essential that we utilize all previous HST imaging to check for variability and optical HSC data to look for a dropout.
@@
[…]
@@ 4 Properties of the TDE Candidate HZTDE-1
Figure 8: Cutouts (top panels) and SED (bottom panel) of high redshift TDE candidate, HZTDE-1. The cutouts include imaging from UltraVISTA, Hyper Suprime Cam, Hubble ACS, and JWST NIRCam. The source is only detected to $3\sigma$ in the NIRCam images. The SED shows the observed SED as red circles, the best fit SLSN spectrum (blue line) and TDE spectrum (red line), and all photometric upper limits as downward-pointing triangles. Errorbars in magnitude exist for the NIRCam data but are too small to be seen. The transmission curve for each filter is plotted below it. A simulated spectrum of TDE 2019qiz at $z=4$ is overplotted as a dotted line.
@@ 4.1 Comparison to supernova models
There are three nearby galaxies that HZTDE-1 could belong to if it is a lower-redshift supernova. We assess its possible association with the galaxy directly above it in the cutouts shown in Figure 8, because it is the closest galaxy, with an angular separation of 0.9980.998”, the brightest galaxy with $M_{F444W}=23.7$ AB mag, and the most extended with effective radius $R_{e}=0.20$” as modeled by GALFIT (Peng et al., 2002). The GALFIT model for the sources in the field is shown in Figure 10, alongside the F [^4] W image and the residual. The best-fit galaxy spectrum to this source’s SED has a photometric redshift $z_{\rm phot}=1.75^{+0.1}_{-0.07}$. This galaxy has a slightly above average best-fit specific star formation rate given its photometric redshift, of $\log(\rm{sSFR})/M_{\odot}~yr^{-1}=-8.67$. An especially high star formation rate would support the occurrence of a supernova in its outskirts (Botticella et al., 2012). The SED is well-sampled from the HSC $g$-band to NIRCam F [^4] W, and the model spectrum fits accurately with a $\chi^{2}=1.72$ per degree of freedom. We follow the technique developed in Gupta et al. (2016) to assess host galaxy association for a given supernova through calculating the directional light radius (DLR), which is the elliptical radius of a galaxy in the direction of the supernova. They then calculate the $d_{\rm DLR}$, which is the distance between the supernova and the galaxy normalized by the DLR. They utilize both $HST$ ACS data and MICECATv2.0 simulations to show that any SN with a galaxy >5>5 DLR away should be assigned “hostless”, and achieve $\sim 95\%$ completeness with a similar cut. We calculate the unitless dDLR for HZTDE-1 to be $5.05$ relative to the nearest galaxy, making it highly likely to be a hostless transient. This dDLR is even higher for the short-wavelength NIRCam filters (>6>6 DLR in F115W), which have higher spatial resolution.
Figure 10: GALFIT modeling of the morphology of the sources in the F [^4] W imaging of HZTDE-1. We display the size of 11 directional light radius (DLR) from the galaxy to the transient to show the accuracy of the modeling. The HZTDE-1 is >5>5 DLR from the galaxy, unlikely to be associated. In shorter-wavelength filters, this DLR separation increases.
[…]
@@
There is a possibility that HZTDE-1 is a superluminous supernova (SLSN), as shown in Figure 8. SLSNe have similar intrinsic luminosities to TDEs and their spectral continua are primarily blackbody thermal emission (Gomez et al., 2024). Their host galaxies tend to be less massive, meaning that at large enough distances ($z\gtrsim 3$) they may also outshine their hosts and be point sources. Therefore, to be certain whether HZTDE-1 is a high-redshift TDE or high-redshift hostless supernova, either follow-up spectroscopy or monitoring of its photometric evolution is necessary. With a single photometric detection, the relative rates of TDEs and SLSNe at $z>3$ is the main factor that determines what HZTDE-1 likely is. Current best estimates of the volumetric rate of SLSNe place them somewhere between $0.91\times 10^{-7}~\rm{SNe}/\rm{year}/\rm{Mpc}^{3}$ at $z=1$ (Prajs et al., 2017) and $1.99\times 10^{-7}~\rm{SNe}/\rm{year}/\rm{Mpc}^{3}$ at $z=0.2$ (Quimby et al., 2013), as compared to the current best estimate for the local optical TDE volumetric rate of $3.1\times 10^{-7}~\rm{TDEs}/\rm{year}/\rm{Mpc}^{3}$. Tanaka et al. (2013) predicts the volumetric rate at SLSNe at higher redshifts (up to $z\geq 10$) by scaling it with the cosmic star formation rate (SFR) density using two SFR models (Robertson & Ellis, 2012; Hopkins & Beacom, 2006). They calibrate to the SLSN low-redshift rate, and account for the initial mass function and the fraction of massive stars to find that the SLSN volumetric rate increases with redshift until $z\sim 5$, where the SFR density peaks, and then the rate decreases with redshift indefinitely. At its maximum, $z\sim 5$, the volumetric SLSN rate is $10\times$ the local rate. We expect the TDE rate to also be elevated by at least $10\times$ the local rate, as shown in Karmen & et al. (in prep.). HZTDE-1 would need to be blue for a SLSN, with a best-fit temperature 24, 000 K at best-fit redshift z = 3.6, but this is possible briefly before peak and during peak brightness. However, this fit was performed without a prior on SLSN behavior and a more reasonable 19,000 K SN will fit at z=3.2. Indeed, because of the low-mass nature of SLSN hosts, the SN could occur in a lower redshift, very faint host galaxy and would be intrinsically redder. Given the sample of host galaxies for the Palomar Transient Factory observed SLSNe (Perley et al., 2016), these hosts could drop out of the COSMOS-Web survey as soon as $z\sim 1.7$. This is around the peak of cosmic star formation rates and would imply a reasonable number of hostless SLSNe. Therefore, without further monitoring, the possibility of a lower-z SLSN is equally likely to a TDE.
[…]
@@ 5 Discussion and Future Confirmation
In this work, we model the properties of TDEs at high redshift assuming that their populations are similar to those at low redshift. We argue that the assumptions made about their host galaxies (Section 2.1) and TDE spectra (Section 2.3) set limits on the colors and morphologies of high-redshift TDEs. However, any number of these assumptions can be relaxed. In future work, for example, modeling of TDEs in extended hosts, or modeling of dust-extincted TDE emission can increase the efficiency of searches for TDEs in wide-field surveys. As this search offers an opportunity to study the UV emission shortward of $2000$ Å which is usually not observed, deviations from blackbody assumptions about TDE spectra, such as those mentioned in Section 2.4 may emerge from studies of high-redshift TDEs.
Either spectroscopic or photometric follow-up could be used to confirm HZTDE-1’s classification as a TDE and its redshift. TDEs have blue spectra with blackbody continua and broad Hydrogen and/or Helium emission features that narrow as the TDE evolves. A key distinction between TDEs and supernovae is their decline: TDEs do not change temperature as they cool (van Velzen et al., 2020), while supernovae cool rapidly after peak, typically 5000-8000 K (Gomez et al., 2024; Aamer et al., 2025). Without imaging of a host galaxy to verify that HZTDE-1 is nuclear, either identification of spectroscopic features or photometric monitoring is needed to confidently confirm it as a TDE. However, because of the degeneracy between redshift, luminosity, and temperature, a spectrum of the TDE or host would be needed to calculate a redshift. Spectroscopy, specifically in the infrared with JWST’s NIRSpec, is key to observing common TDE emission lines. X-ray emission is not expected because, as mentioned, the CGM will absorb all of the soft X-ray emission characteristic of TDEs at $z>3$. Longer wavelengths may be worth monitoring for later-time infrared flares due to dust echoes (van Velzen et al., 2016).
If HZTDE-1 is a high-redshift SLSN, its discovery and classification will be valuable for understanding stellar evolution at higher redshift. SLSNe are expected to result from the explosions of massive stars (up to tens of solar masses Blanchard et al., 2020) and have a strong preference for galaxies with low metallicities (Perley et al., 2016; Schulze et al., 2018; Cleland et al., 2023). To first order, the cosmic rates of SLSNe should follow the cosmic star formation rate (Tanaka et al., 2013). Deviations from this relationship could provide insight into the abundance of metal-poor galaxies at earlier cosmic times. The methodology outlined in this paper could identify potentially hundreds of high-redshift SLSNe in upcoming wide-field surveys (e.g. the Roman High Latitude Wide Area Survey) and therefore could lead to measurements of how their rates and hosts change with redshift. This in turn can be used to study how massive stars form and evolve as the metallicity of the universe changes over cosmic time.
@@ 5.1 TDE rates in COSMOS-Web
In the companion work Karmen & et al. (in prep.), we calculate observed TDE rates as a function of redshift in LSST, Roman, and the COSMOS-Web survey. We find that if we simply extrapolate the local TDE rate, scaled by the number density of SMBHs (Shankar et al., 2009) that can disrupt a main-sequence star (following Kochanek, 2016; Sun et al., 2015; Donnarumma et al., 2015), we get a $20\%$ chance of finding a single TDE in the COSMOS-Web survey.
We show, however, that the observed properties of high-redshift galaxies would imply a dramatically enhanced rate of TDEs at high-redshift $z>5$, and perhaps as low as $z=2$. Namely, many galaxies are observed to be centrally compact (Finkelstein et al., 2023; Kartaltepe et al., 2023; Baggen et al., 2023; Guia et al., 2024), and may host nuclear stellar clusters (Ricotti et al., 2016; Sun et al., 2023) at high redshifts. This increases central stellar density, $\rho_{\star}$, by 121-2 orders of magnitude (Baggen et al., 2024) which in turn can easily lead to a $\times 100$ TDE rate enhancement assuming a galaxy’s TDE rate scales as $\propto\rho_{\star}^{0.65}$ (Pfister et al., 2020), or even higher given the more dramatic observed scaling with stellar density $\propto\rho_{\star}^{0.9}$ (Graur et al., 2018). This enhancement is in agreement with simulations of TDEs in galaxies with nuclear stellar clusters (Kritos et al., 2024; Pfister et al., 2020). Additionally, galaxy mergers are shown to enhance TDE rates both observationally and theoretically. Host studies of optical/UV TDEs demonstrate a clear preference for galaxies which are believed to be products of mergers (French et al., 2020; Wevers & French, 2024), and the TDE rate is seen in hydrodynamical simulations to be enhanced by an order of magnitude for 100 Myr after a merger (Pfister et al., 2021). With a galaxy merger rate that increases as a function of redshift, peaking at cosmic noon (Ventou et al., 2017; Duncan et al., 2019; Perna et al., 2025) we predict TDE rates to be enhanced from mergers as well. In previous works (e.g. Kochanek, 2016; Donnarumma et al., 2015), the volumetric TDE rate has been scaled with a quickly-evolving SMBH mass function, as as a result it was expected that TDE rates decrease rapidly with the SMBH mass function as a function of redshift. However, given the abundance of supermassive and overmassive black holes detected at early cosmic times (e.g. Maiolino et al., 2023; Larson et al., 2023; Kokorev et al., 2023), this phenomenon will be significantly reduced. TDEs are among the few luminous phenomena very close to an SMBH, and are believed to produce super-Eddington accretion at early times (Strubbe & Quataert, 2009) so their study at high redshifts can give an indication to the mechanisms by which early SMBHs accrete and grow.